I'm getting stuck at the end of this one...any help would be greatly appreciated!a pen with one side against a wall WITH 6M of fencing. X is the width and the length is (6-2x). The area is a=x(6-2x).
A). What dimensions of the pen will produce the maximum area?
B) what is the maximum area?
I have done the following
A=(6x-2x)
Max area occurs at axis of symmetry
X=-6/-4
3/2 is the max area
L=6-2(3/2)
Length=5.5?
I have a hard time working with fractions...any help would be greatly appreciated!

Answer by ankor@dixie-net.com(22740)

a pen with one side against a wall WITH 6M of fencing. X is the width and the length is (6-2x). The area is a=x(6-2x).
A). What dimensions of the pen will produce the maximum area?
B) what is the maximum area?
I have done the following
right here it should be
A=(6x-2x^2)
Max area occurs at axis of symmetry
X=-6/-4
x = +1.5, we can use decimals here
:
Find the Area by substituting for x in the Area equation
A = 6(1.5) - 2(1.5^2)
A = 9 - 2(2.25
A - 9 - 4.5
A = 4.5 sq/m is the max area
:
:
Check this using the dimensions
L = 6 - 2(1.5)
L = 3
W = 1.5
3 * 1.5 = 4.5